# CatCMA with Margin: Stochastic Optimization for Continuous, Integer, and Categorical Variables

### v5

@@ 4. Proposed Integer Handling
In this section, we propose the new integer handling suitable for integrating into CatCMA based on the previous section.
@@ 4.1. Modified Margin Correction
(33) $\exists i\in\{1\!:\!\lambda,\ldots,\mu\!:\!\lambda\}\enspace\text{s.t.}\enspace[\boldsymbol{z}_{i}]_{n}\neq[\textsc{Enc}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$
@@
Case of ${}[\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,1},z_{n,L_{n}}\}$.
@@
Case of ${}[\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,2},\ldots,z_{n,L_{n}-1}\}$.
The updates of (18) to (19) are modified to the following:
@@ 4.2. Introduction of Integer Centering
After this correction, $[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ is corrected using $p^{({t+1})}_{\mathrm{mut},{n}}$ previously calculated in (35) as follows:
When $\langle\boldsymbol{A}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}$ is corrected to the lower bound in (39) and $p^{({t+1})}_{\mathrm{mut},{n}}$ is $\alpha$, $[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ is corrected to match $\boldsymbol{z}_{n,1}$ or $\boldsymbol{z}_{n,L_{n}}$, as shown in the bottom of Figure 1. A summary of the modified margin correction is provided in the supplementary material to facilitate following the revised and added update rules.
@@ 4.3. Derivation of Promising Margin Settings
Before the derivation of promising margin settings, we confirm that the modified margin correction can suppress unnecessary fluctuations of the mutation rate. The last two columns of Figure 2 show the results of a typical trial introducing the modified margin correction and the integer centering. Since no upper bound is imposed on the mutation rate in the case of a successful integer mutation, the impact on integer variable optimization is minimal. Moreover, when the optimization of integer variables has been sufficiently performed, a successful integer mutation does not occur, and the upper bound prevents the mutation rate from fluctuating. Therefore, we can reasonably assume the convergence of the mutation rate for deriving promising margin settings.
@@
(41) $[\textnormal{{Enc}}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$ $=\boldsymbol{z}^{\mathrm{opt}}_{n}\enspace\text{for all}\enspace n\in\{1,\ldots,{N_{\mathrm{in}}}\}$
(42) $p^{({t})}_{\mathrm{mut},{n}}$ $=\alpha\enspace\text{for all}\enspace n\in\{1,\ldots,{N_{\mathrm{in}}}\}$
@@
(43) $\alpha=1-0.73^{\frac{1}{{N_{\mathrm{in}}}}}\enspace,$
@@
(44) $\Pr\left(\lambda_{\mathrm{non}}\leq\lambda-\left\lfloor\frac{\lambda}{2}\right\rfloor\right)\geq 0.95\enspace.$
The proof of Proposition 1 can be found in the supplementary material. This proposition provides the margin that ensures that the top $\mu=\lfloor\lambda/2\rfloor$ solutions do not contain non-optimal integers with high probability when the part of the multivariate Gaussian distribution responsible for generating integer variables has converged to the optimal state. This implies that, in the updates of (7) to (9), the solutions containing non-optimal integers are almost ignored. On the other hand, when using negative weights, the update of the covariance matrix (11) is affected by such samples, which will be investigated in the next subsection.
@@ 4.4. Experimental Evaluation of Proposed Margin Correction
In EllipsoidInt, integer variables have a greater impact on the evaluation value than continuous variables, whereas in REllipsoidInt, continuous variables have a greater impact than integer variables. The numbers of dimensions for both continuous and integer variables were set to $20$. The integer space was set as $\mathcal{Z}_{1}=\cdots=\mathcal{Z}_{20}=\{-10,-9,\ldots,9,10\}$. The initial parameters of the multivariate Gaussian distribution were given as $\boldsymbol{C}^{(0)}=\mathbf{I}_{{N_{\mathrm{mi}}}}$ and $\sigma^{(0)}=1$, and $\boldsymbol{m}^{(0)}$ was uniformly sampled in $[1,3]^{{N_{\mathrm{mi}}}}$. In the LB+IC-CMA-ES, we used the settings of IC-LBx1 and IC-LBx2 in (Marty et al., 2024). The other settings were the same as those in Section 3.2.
@@ Results.
0:  $\boldsymbol{m}^{(0)},\sigma^{(0)},\boldsymbol{C}^{(0)},\boldsymbol{q}^{(0)}$
1:  $\boldsymbol{A}^{(0)}=\mathbf{I}_{N_{\mathrm{mi}}}$, $\boldsymbol{p}^{(0)}_{\sigma}=\boldsymbol{p}^{(0)}_{c}=\boldsymbol{0}$, $\delta^{(0)}=1$, $\boldsymbol{s}^{(0)}=\boldsymbol{0}$, $\gamma^{(0)}=0$, $p^{({0})}_{\mathrm{mut},{1}}=\cdots=p^{({0})}_{\mathrm{mut},{{N_{\mathrm{in}}}}}=1$, $t=0$
@@
9:           $[\boldsymbol{x}_{i}]_{n}\leftarrow[\textsc{Enc}(\boldsymbol{v}_{i})]_{j^{\mathrm{co}}_{n}}$
@@
12:           $[\boldsymbol{z}_{i}]_{n}\leftarrow[\textsc{Enc}(\boldsymbol{v}_{i})]_{j^{\mathrm{in}}_{n}}$
@@
15:     Evaluate $(\boldsymbol{x}_{1},\boldsymbol{z}_{1},\boldsymbol{c}_{1}),\ldots,(\boldsymbol{x}_{\lambda},\boldsymbol{z}_{\lambda},\boldsymbol{c}_{\lambda})$ on $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})$.
16:     Correct $\boldsymbol{v}_{1:\lambda},\ldots,\boldsymbol{v}_{\mu:\lambda}$ and $\boldsymbol{y}_{1:\lambda},\ldots,\boldsymbol{y}_{\mu:\lambda}$ by integer centering in (Marty et al., 2024).
17:     Update $\boldsymbol{m}^{(t)}$, $\boldsymbol{p}^{(t)}_{\sigma}$, $\boldsymbol{p}^{(t)}_{c}$, $\boldsymbol{C}^{(t)}$, and $\sigma^{(t)}$ according to Section 2.2.1.
18:     Update $\boldsymbol{q}^{(t)}$, $\boldsymbol{s}^{(t)}$, $\gamma^{(t)}$, $\delta^{(t)}$, and modify $\sigma^{(t+1)}$ and $\boldsymbol{q}^{(t+1)}$ according to Section 2.3.
19:     Update $\boldsymbol{m}^{(t+1)}$, $\boldsymbol{A}^{(t)}$, $p^{({t})}_{\mathrm{mut},{1}},\ldots,p^{({t})}_{\mathrm{mut},{{N_{\mathrm{in}}}}}$ by the modified margin correction.
@@
Table 1. Benchmark functions to be minimized.
Name Definition
SphereIntCOM $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}}\boldsymbol{x}_{n}^{2}+\sum_{n=1}^{{N_{\mathrm{in}}}}\boldsymbol{z}_{n}^{2}+{N_{\mathrm{ca}}}-\sum_{n=1}^{{N_{\mathrm{ca}}}}\boldsymbol{c}_{n,1}$
EllipsoidIntCLO $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}}10^{6\frac{n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{x}_{n}^{2}+\sum_{n=1}^{{N_{\mathrm{in}}}}10^{6\frac{{N_{\mathrm{co}}}+n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{z}_{n}^{2}+{N_{\mathrm{ca}}}-\sum_{n=1}^{{N_{\mathrm{ca}}}}\prod_{n^{\prime}=1}^{n}\boldsymbol{c}_{n^{\prime},1}$
REllipsoidIntCLO $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}}10^{6\frac{{N_{\mathrm{in}}}+n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{x}_{n}^{2}+\sum_{n=1}^{{N_{\mathrm{in}}}}10^{6\frac{n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{z}_{n}^{2}+{N_{\mathrm{ca}}}-\sum_{n=1}^{{N_{\mathrm{ca}}}}\prod_{n^{\prime}=1}^{n}\boldsymbol{c}_{n^{\prime},1}$
MVProximity $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}(={N_{\mathrm{ca}}})}\bigl{(}\frac{\boldsymbol{x}_{n}}{x_{\max}}-\boldsymbol{\zeta}_{n}\bigr{)}^{2}+\sum_{n=1}^{{N_{\mathrm{in}}}(={N_{\mathrm{ca}}})}\bigl{(}\frac{\boldsymbol{z}_{n}}{z_{\max}}-\boldsymbol{\zeta}_{n}\bigr{)}^{2}+\sum_{n=1}^{{N_{\mathrm{ca}}}}\boldsymbol{\zeta}_{n}$  ,  where $\boldsymbol{\zeta}_{n}=\frac{1}{K_{n}}\sum_{k=1}^{K_{n}}(k-1)\boldsymbol{c}_{n,k}$  .
@@ 5. Proposed Method : CatCMA with Margin
(46) $[\textnormal{{Enc}}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$ $=\boldsymbol{z}^{\mathrm{opt}}_{n}\enspace\text{for all}\enspace n\in\{1,\ldots,{N_{\mathrm{in}}}\}$
(47) $p^{({t})}_{\mathrm{mut},{n}}$ $=\alpha\enspace\text{for all}\enspace n\in\{1,\ldots,{N_{\mathrm{in}}}\}$
Without loss of generality, assume that categories of the optimal solution are the first categories in all dimensions and the parameter of the categorical distribution satisfies
(48) $\begin{split}&\boldsymbol{q}^{(t)}_{n,1}=1-\boldsymbol{q}^{\mathrm{min}}_{n}(K_{n}-1)\enspace,\\ &\boldsymbol{q}^{(t)}_{n,k}=\boldsymbol{q}^{\mathrm{min}}_{n}\quad\text{for all}\enspace k\in\{2,\ldots,K_{n}\}\end{split}$
@@
(49) $\alpha$ $=1-0.73^{\frac{1}{{N_{\mathrm{in}}}+{N_{\mathrm{ca}}}}}\enspace,$
(50) $\boldsymbol{q}^{\mathrm{min}}_{n}$ $=\frac{1}{K_{n}-1}\left(1-0.73^{\frac{1}{{N_{\mathrm{in}}}+{N_{\mathrm{ca}}}}}\right)\enspace\text{for all}\enspace n\in\{1,\ldots,{N_{\mathrm{ca}}}\}\enspace,$
@@
(51) $\Pr\left(\lambda_{\mathrm{non}}\leq\lambda-\left\lfloor\frac{\lambda}{2}\right\rfloor\right)\geq 0.95\enspace.$
The proof of Proposition 2 can be found in the supplementary material. This proposition focuses on the phase where optimization of both integer and categorical variables has been sufficiently performed. This margin setting is expected to achieve a balance between avoiding the fixation of discrete variables and efficient continuous variable optimization.
@@
6. Experiments
This section evaluates the search performance of CatCMAwM on MV-BBO problems. We used the following benchmark functions.
@@ 6. Experiments
These definitions are summarized in Table 1. The first three functions are the sum of existing mixed-integer and categorical benchmark functions. In MVProximity, there are dependencies between continuous and categorical variables, and between integer and categorical variables.
Experimental Settings.
The integer space was set as $\mathcal{Z}_{1}=\cdots=\mathcal{Z}_{{N_{\mathrm{in}}}}=\{-3,-2,\ldots,2,3\}$. The numbers of categories were set to $5$. In MVProximity, we set $x_{\smash{\max}}$ and $z_{\smash{\max}}$ to $3$ according to the search space. As baseline state-of-the-art methods, we experimented with CASMOPOLITAN (Wan et al., 2021), SMAC3 (Hutter et al., 2011; Lindauer et al., 2022), and Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) implemented by Optuna (Akiba et al., 2019). In the three methods, the ranges of the continuous variables were set to $[-3,3]$. The evaluation budget on the CASMOPOLITAN was set to $400$ due to long internal computation time. To confirm the effectiveness of the proposed integer handling, we also experimented with two CatCMA methods naively incorporating the default margin correction and the integer handling of LB+IC-CMA-ES, using the default hyperparameters such as the margin value for each method. In the CatCMA-based methods, the initial distribution parameters were given as $\boldsymbol{C}^{(0)}=\mathbf{I}_{{N_{\mathrm{mi}}}}$ and $\sigma^{(0)}=1$, $\boldsymbol{q}_{n}^{\smash{(0)}}=1/K_{n}$ and $\boldsymbol{m}^{(0)}$ was uniformly sampled in $[1,3]^{\smash{{N_{\mathrm{mi}}}}}$, and sample size was set as $\lambda=4+\lfloor 3\ln({N_{\mathrm{co}}}+{N_{\mathrm{in}}}+{N_{\mathrm{ca}}})\rfloor$.
Figure 4. The medians and inter-quartile ranges of the best evaluation values for $({N_{\mathrm{co}}},{N_{\mathrm{in}}},{N_{\mathrm{ca}}})=(4,4,4)$ over 20 independent trials.

### v6

@@ 4. Proposed Integer Handling
In this section, we propose a new integer handling suitable for integrating into CatCMA based on the previous section.
@@ 4.1. Modified Margin Correction
(33) $\exists i\in\{1\!:\!\lambda,\ldots,\mu\!:\!\lambda\}\kern 5.0pt\text{s.t.}\kern 5.0pt[\boldsymbol{z}_{i}]_{n}\neq[\textsc{Enc}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$
@@
Case of $\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,1},z_{n,L_{n}}\}$.
@@
Case of $\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,2},\ldots,z_{n,L_{n}-1}\}$.
The updates (18) to (19) are modified to the following:
@@ 4.2. Introduction of Integer Centering
After this correction, $\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ is corrected using $p^{({t+1})}_{\mathrm{mut},{n}}$ previously calculated in (35) as follows:
(40) $\begin{split}&[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}\leftarrow{\ell^{({t+1})}_{\mathrm{neigh},{n}}}+\mathrm{sign}\left([\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}-{\ell^{({t+1})}_{\mathrm{neigh},{n}}}\right)\\ &\qquad\qquad\qquad\qquad\qquad\cdot\sigma^{(t+1)}\langle\boldsymbol{A}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}\sqrt{\langle\boldsymbol{C}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{mut},{n}}})}\end{split}$
When $\langle\boldsymbol{A}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}$ is corrected to the lower bound in (39) and $p^{({t+1})}_{\mathrm{mut},{n}}$ is $\alpha$, $\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ is corrected to match $\boldsymbol{z}_{n,1}$ or $\boldsymbol{z}_{n,L_{n}}$, as shown in the bottom of Figure 1. A summary of the modified margin correction is provided in Appendix A to facilitate following the revised and added update rules.
@@ 4.3. Derivation of Promising Margin Settings
Before the derivation of promising margin settings, we confirm that the modified margin correction can suppress unnecessary fluctuations of the mutation rate. The results of a typical trial introducing the modified margin correction and the integer centering are included in Figure 2. Since no upper bound is imposed on the mutation rate in the case of a successful integer mutation, the impact on integer variable optimization is minimal. Moreover, when the optimization of integer variables has been sufficiently performed, a successful integer mutation does not occur, and the upper bound prevents the mutation rate from fluctuating. Therefore, we can reasonably assume the convergence of the mutation rate for deriving promising margin settings.
@@
(41) $[\textnormal{{Enc}}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$ $=\boldsymbol{z}^{\mathrm{opt}}_{n}\kern 5.0pt\text{for all}\kern 5.0ptn\in\{1,\ldots,{N_{\mathrm{in}}}\}$
(42) $p^{({t})}_{\mathrm{mut},{n}}$ $=\alpha\kern 5.0pt\text{for all}\kern 5.0ptn\in\{1,\ldots,{N_{\mathrm{in}}}\}$
@@
(43) $\alpha=1-0.73^{\frac{1}{{N_{\mathrm{in}}}}}\kern 5.0pt,$
@@
(44) $\Pr\left(\lambda_{\mathrm{non}}\leq\lambda-\left\lfloor\frac{\lambda}{2}\right\rfloor\right)\geq 0.95\kern 5.0pt.$
The proof of Proposition 1 can be found in Appendix B. This proposition provides the margin that ensures that the top $\mu=\lfloor\lambda/2\rfloor$ solutions do not contain non-optimal integers with high probability when the part of the multivariate Gaussian distribution responsible for generating integer variables has converged to the optimal state. This implies that, in the updates (7) to (9), the solutions containing non-optimal integers are almost ignored. On the other hand, when using negative weights, the update of the covariance matrix (11) is affected by such samples, which will be investigated in the next subsection.
@@ 4.4. Experimental Evaluation of Proposed Margin Correction
In EllipsoidInt, integer variables have a greater impact on the evaluation value than continuous variables, whereas in REllipsoidInt, continuous variables have a greater impact than integer variables. The numbers of dimensions for both continuous and integer variables were set to $20$. The integer space was set as $\mathcal{Z}_{1}=\cdots=\mathcal{Z}_{20}=\{-10,-9,\ldots,9,10\}$. The initial parameters of the multivariate Gaussian distribution were given as $\boldsymbol{C}^{(0)}=\mathbf{I}_{{N_{\mathrm{mi}}}}$ and $\sigma^{(0)}=1$, and $\boldsymbol{m}^{(0)}$ was uniformly sampled in $1,3]^{{N_{\mathrm{mi}}}}$. In the LB+IC-CMA-ES, we used the settings of IC-LBx1 and IC-LBx2 in (Marty et al., 2024). The other settings were the same as those in Section 3.2.
@@ Results.
[…]
@@
9:    $\boldsymbol{x}_{i}]_{n}\leftarrow[\textsc{Enc}(\boldsymbol{v}_{i})]_{j^{\mathrm{co}}_{n}}$
@@
12:    $\boldsymbol{z}_{i}]_{n}\leftarrow[\textsc{Enc}(\boldsymbol{v}_{i})]_{j^{\mathrm{in}}_{n}}$
@@
15:  Evaluate $(\boldsymbol{x}_{1},\boldsymbol{z}_{1},\boldsymbol{c}_{1}),\ldots,(\boldsymbol{x}_{\lambda},\boldsymbol{z}_{\lambda},\boldsymbol{c}_{\lambda})$ on $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})$
16:  Correct $\boldsymbol{v}_{1:\lambda},\ldots,\boldsymbol{v}_{\mu:\lambda}$ and $\boldsymbol{y}_{1:\lambda},\ldots,\boldsymbol{y}_{\mu:\lambda}$ by integer centering in (Marty et al., 2024)
17:  Update $\boldsymbol{m}^{(t)}$, $\boldsymbol{p}^{(t)}_{\sigma}$, $\boldsymbol{p}^{(t)}_{c}$, $\boldsymbol{C}^{(t)}$, and $\sigma^{(t)}$ according to Section 2.2.1
18:  Update $\boldsymbol{q}^{(t)}$, $\boldsymbol{s}^{(t)}$, $\gamma^{(t)}$, $\delta^{(t)}$, and modify $\sigma^{(t+1)}$ and $\boldsymbol{q}^{(t+1)}$ according to Section 2.3
19:  Update $\boldsymbol{m}^{(t+1)}$, $\boldsymbol{A}^{(t)}$, $p^{({t})}_{\mathrm{mut},{1}},\ldots,p^{({t})}_{\mathrm{mut},{{N_{\mathrm{in}}}}}$ by the modified margin correction
@@ 5. Proposed Method: CatCMA with Margin
(46) $[\textnormal{{Enc}}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$ $=\boldsymbol{z}^{\mathrm{opt}}_{n}\kern 5.0pt\text{for all}\kern 5.0ptn\in\{1,\ldots,{N_{\mathrm{in}}}\}$
(47) $p^{({t})}_{\mathrm{mut},{n}}$ $=\alpha\kern 5.0pt\text{for all}\kern 5.0ptn\in\{1,\ldots,{N_{\mathrm{in}}}\}$
Without loss of generality, assume that categories of the optimal solution are the first categories in all dimensions. Assume that the parameter of the categorical distribution satisfies
(48) $\begin{split}&\boldsymbol{q}^{(t)}_{n,1}=1-\boldsymbol{q}^{\mathrm{min}}_{n}(K_{n}-1)\kern 5.0pt,\\ &\boldsymbol{q}^{(t)}_{n,k}=\boldsymbol{q}^{\mathrm{min}}_{n}\quad\text{for all}\kern 5.0ptk\in\{2,\ldots,K_{n}\}\end{split}$
@@
(49) $\alpha$ $=1-0.73^{\frac{1}{{N_{\mathrm{in}}}+{N_{\mathrm{ca}}}}}\kern 5.0pt,$
(50) $\boldsymbol{q}^{\mathrm{min}}_{n}$ $=\frac{1}{K_{n}-1}\left(1-0.73^{\frac{1}{{N_{\mathrm{in}}}+{N_{\mathrm{ca}}}}}\right)\kern 5.0pt\text{for all}\kern 5.0ptn\in\{1,\ldots,{N_{\mathrm{ca}}}\}\kern 5.0pt,$
@@
(51) $\Pr\left(\lambda_{\mathrm{non}}\leq\lambda-\left\lfloor\frac{\lambda}{2}\right\rfloor\right)\geq 0.95\kern 5.0pt.$
The proof of Proposition 2 can be found in Appendix C. This proposition focuses on the phase where optimization of both integer and categorical variables has been sufficiently performed. This margin setting is expected to achieve a balance between avoiding the fixation of discrete variables and efficient continuous variable optimization.
@@
6. Experiments on Single-Objective Optimization Problems
This section evaluates the search performance of CatCMAwM on single-objective MV-BBO problems. We used the following benchmark functions.
@@ 6. Experiments on Single-Objective Optimization Problems
    (52) $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}}\boldsymbol{x}_{n}^{2}+\sum_{n=1}^{{N_{\mathrm{in}}}}\boldsymbol{z}_{n}^{2}+{N_{\mathrm{ca}}}-\sum_{n=1}^{{N_{\mathrm{ca}}}}\boldsymbol{c}_{n,1}$
@@
    (53) $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}}10^{6\frac{n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{x}_{n}^{2}+\sum_{n=1}^{{N_{\mathrm{in}}}}10^{6\frac{{N_{\mathrm{co}}}+n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{z}_{n}^{2}+{N_{\mathrm{ca}}}-\sum_{n=1}^{{N_{\mathrm{ca}}}}\prod_{n^{\prime}=1}^{n}\boldsymbol{c}_{n^{\prime},1}$
@@
    (54) $f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}}10^{6\frac{{N_{\mathrm{in}}}+n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{x}_{n}^{2}+\sum_{n=1}^{{N_{\mathrm{in}}}}10^{6\frac{n-1}{{N_{\mathrm{co}}}+{N_{\mathrm{in}}}-1}}\boldsymbol{z}_{n}^{2}+{N_{\mathrm{ca}}}-\sum_{n=1}^{{N_{\mathrm{ca}}}}\prod_{n^{\prime}=1}^{n}\boldsymbol{c}_{n^{\prime},1}$
@@
    (55) $\begin{split}f(\boldsymbol{x},\boldsymbol{z},\boldsymbol{c})=\sum_{n=1}^{{N_{\mathrm{co}}}(={N_{\mathrm{ca}}})}\left(\frac{\boldsymbol{x}_{n}}{x_{\mathrm{scale}}}-\boldsymbol{\zeta}_{n}\right)^{2}&+\sum_{n=1}^{{N_{\mathrm{in}}}(={N_{\mathrm{ca}}})}\left(\frac{\boldsymbol{z}_{n}}{z_{\mathrm{scale}}}-\boldsymbol{\zeta}_{n}\right)^{2}+\sum_{n=1}^{{N_{\mathrm{ca}}}}\boldsymbol{\zeta}_{n}\kern 5.0pt,\\ &\qquad\qquad\text{where}\kern 5.0pt\boldsymbol{\zeta}_{n}=\frac{1}{K_{n}}\sum_{k=1}^{K_{n}}(k-1)\boldsymbol{c}_{n,k}\kern 5.0pt.\end{split}$
The first three functions are the sum of existing mixed-integer and categorical benchmark functions. In MVProximity, there are dependencies between continuous and categorical variables, and between integer and categorical variables.
Experimental Settings
The integer space was set as $\mathcal{Z}_{1}=\cdots=\mathcal{Z}_{{N_{\mathrm{in}}}}=\{-3,-2,\ldots,2,3\}$. The numbers of categories were set to $5$. In MVProximity, we set $x_{\smash{\mathrm{scale}}}$ and $z_{\smash{\mathrm{scale}}}$ to $3$ according to the search space. As baseline state-of-the-art methods, we experimented with CASMOPOLITAN (Wan et al., 2021), SMAC3 (Hutter et al., 2011; Lindauer et al., 2022), and Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) implemented by Optuna (Akiba et al., 2019). In the three methods, the ranges of the continuous variables were set to [3,3][-3,3]. The evaluation budget on the CASMOPOLITAN was set to $400$ due to long internal computation time. To confirm the effectiveness of the proposed integer handling, we also experimented with two CatCMA methods naively incorporating the default margin correction and the integer handling of LB+IC-CMA-ES, using the default hyperparameters such as the margin value for each method. In the CatCMA-based methods, the initial distribution parameters were given as $\boldsymbol{C}^{(0)}=\mathbf{I}_{{N_{\mathrm{mi}}}}$ and $\sigma^{(0)}=1$, $\boldsymbol{q}_{n}^{\smash{(0)}}=1/K_{n}$ and $\boldsymbol{m}^{(0)}$ was uniformly sampled in $1,3]^{\smash{{N_{\mathrm{mi}}}}}$, and sample size was set as $\lambda=4+\lfloor 3\ln({N_{\mathrm{co}}}+{N_{\mathrm{in}}}+{N_{\mathrm{ca}}})\rfloor$.